OpenStudy (anonymous):

Solve the system of equations by graphing. 3x-y=3, 4x+y=4

OpenStudy (anonymous):

mark x=1 and y=-3 ------(1) mark x=1 and y=-4 ------(1) and plot the graph :D

OpenStudy (anonymous):

did you understand @marixoxo76 ?

OpenStudy (anonymous):

i need two plots for each equation :/

OpenStudy (jdoe0001):

so, graph them both, where they INTERSECT each other, that's where the X and Y have common values, or are equal, and thus is the answer

OpenStudy (radar):

Locate each equation intercepts (x and y) The first equation let y=0 getting x=3, then let x=0 getting y=-3 Those are the two intercepts for first equation (3,0) and (0,-3) plot that line. Do you need further help for the second equation?

OpenStudy (anonymous):

yes plese @radar

OpenStudy (radar):

O.k I will walk you through this, we are goiing to locate the x and y intercepts for the second equation. First let y = 0 then the equation becomes 4x + 0 = 4 then 4x=4 so x = 4/4=1. That fixes the first point at (1,0) One point on the second line is (1,0) Did that make any sense??

OpenStudy (radar):

Now to find the second point, let x=0 and solve for y (2nd equation)) 4x + y = 4 becomes 0 + y =4 or y=4 second point is (0,4) Your two points to graph the second line are: (0,4) and (1.0) Now get some graph paper and graph those two lines and @jdoe0001 has pointed out where they intersect is the solution values of x and y.

OpenStudy (radar):

In my humble opinion locating those two points on each line is the easiest way to draw the line.

OpenStudy (radar):

The point where they intersect is (1,0) solving by elimination method

OpenStudy (radar):

After establishing the two points for each line, use a straight edge to construct the line, insure you continue the line beyond the points.

OpenStudy (radar):